# FineMatch: Aspect-based Fine-grained Image and Text Mismatch Detection and Correction

### v1

@@ Abstract
1 Introduction
Pretrained vision-language models, such as GPT-4V [1], LLaVA [50], MiniGPT-4 [6], and BLIP [22, 21], have demonstrated impressive ability to perform complex reasoning. Benefiting from large pretrained VLMs, a series of VLM-based methods [13, 23, 53, 16] have emerged and achieved remarkable results on various vision-language (VL) tasks. However, contemporary state-of-the-art VLMs still struggle with fine-grained compositional information understanding [59, 30]. Prior works have pointed out that pretrained VLMs face challenges in comprehending fine-grained visual and textual compositional information [60]. This issue poses a significant limitation to the reliability and performance of VLMs.
In recent years, there has been a growing focus on evaluating and improving the compositionality in large VL models [8, 35, 36]. Most of these approaches focus on constructing hard negative text-image pairs to evaluate the models’ compositionality [64, 49, 55, 59, 12, 31]. These evaluations typically require models to identify the hard negative samples at the sentence level but ignore evaluating the model’s capability to localize the mismatched phrases and provide the corresponding corrections. Nevertheless, evaluating the compositionality of pretrained VL models from the perspective of sentence level may seem almost too trivial a task [8]. Additionally, research into the ability to identify and rectify discrepancies between images and text has been overlooked. Based on this background, we propose FineMatch, a new challenging benchmark for boosting VLMs’ ability to identify and address the fine-grained discrepancies between visual and textual data.
With FineMatch, we analyze and address the semantic discrepancies between visual and textual data from four aspects: Entity, Relation, Attribute, Number. And we provide examples for these four aspects in Appendix 1. We build FineMatch data from both the image side and text side, aggregating data from multiple sources. The FineMatch benchmark comprises 54,800 high-quality, human-annotated image-text pairs, distributed as 47,800 in the training set, 1,000 in the validation set, and 5,000 in the test set. This data originates from various sources, including: GPT-Synthesized Text Data, Retrieved Image and Text Data, and Diffusion Model-Generated Image Data. Each image-text pair encompasses a varying number of mismatched aspects, ranging from 0 to 3. The collection methods for each data source are described in Section 3.
@@ 3.5 Human Annotation
To standardize the labeling scheme across different data sources and ensure data quality while eliminating potentially harmful content and addressing ethical concerns, we employ a consistent annotation team composed of the same group of workers. The annotation interfaces of different sources of data are shown in Appendix 7.
@@ 3.7 Evaluation
where the , and is the weight of the , and , respectively. In this study, we set the weights , , and . To compute the IoU of the aspect representation, we set a threshold to match the predictions with ground truth, if then the predicted triplets matches the ground truth. For each aspect representation prediction, we compute the final score as:
@@
We also provide the pseudocode 1 for the ITM-IoU calculation.
@@ 3.9 Qualitative Analysis
We present the analysis of human-annotated and GPT-Synthesized data, focusing on the Vera Score Gap, Grammar Score Gap, and CLIP Score Gap. These metrics compare the quality changes of human-annotated mismatched captions against the original captions. Previous research [12] analyzes the artifact bias through the Vera Score Gap and Grammar Score Gap, and employs the adversarial refinement to fix this bias. In this research, we utilize human annotation to fix this bias and add the CLIP Score Gap to quantify the semantic changes of the human-annotated mismatched captions. The score gap is typically calculated as , where the and are the Vera/Grammar/CLIP scores of the initial captions and the human-annotated GPT-Synthesised data. The findings, illustrated in Figure 6, reveal that score gap distribution lies on the positive spectrum, indicating that hard negative samples in the GPT-Synthesized data exhibit a higher likelihood of being nonsensical. The results also reflect a marginally reduced similarity to the initial matching captions. While these effects are within an acceptable range. Moreover, GPT rewrites do not significantly impact grammar fluency or introduce errors compared to the original captions. We also provide more quality analysis from different perspectives about other data sources in Appendix 2.
@@ 5 FineMatch for Text-to-Image Generation Hallucination Detection and Correction
Hallucination issues are prevalent in text-to-image (T2I) generation models, particularly for the text prompts that describe detailed scenes and intricate relationships [41, 25]. Many approaches are proposed for reducing hallucination for T2I models [24, 51, 57]. In this study, we illustrate how FineMatch can help detect and address the mismatch between the text prompts and the generated images for T2I generation models.

### v2

@@ Abstract
Keywords:
Pre-trained Vision-Language Models Aspect-based Image and Text Analysis Compositionality1 Introduction
Pretrained vision-language models, such as GPT-4V [1], LLaVA [55], MiniGPT-4 [7], and BLIP [25, 24], have demonstrated impressive ability to perform complex reasoning. Benefiting from large pretrained VLMs, a series of VLM-based methods [15, 26, 52, 18, 63] have emerged and achieved remarkable results on various vision-language (VL) tasks. However, contemporary state-of-the-art VLMs still struggle with fine-grained compositional information understanding [64, 20]. Prior works have pointed out that pretrained VLMs face challenges in comprehending fine-grained visual and textual compositional information [9, 64]. This issue poses a significant limitation to the reliability and performance of VLMs.
In recent years, there has been a growing focus on evaluating and improving the compositionality in large VL models [9, 37, 38]. Most of these approaches focus on constructing hard negative text-image pairs to evaluate the models’ compositionality [68, 54, 59, 13, 64, 14, 33]. These evaluations typically require models to identify the hard negative samples at the sentence level but ignore evaluating the model’s capability to localize the mismatched phrases and provide the corresponding corrections. Nevertheless, evaluating the compositionality of pretrained VL models from the perspective of sentence level may seem almost too trivial a task [9]. Additionally, research into the ability to identify and rectify discrepancies between images and text has been overlooked. Based on this background, we propose FineMatch, a new challenging benchmark for boosting VLMs’ ability to identify and address the fine-grained discrepancies between visual and textual data.
With FineMatch, we analyze and address the semantic discrepancies between visual and textual data from four aspects: Entity, Relation, Attribute, Number. And we provide examples for these four aspects in supplementary material. We build FineMatch data from both the image side and text side, aggregating data from multiple sources. The FineMatch benchmark comprises 49,906 high-quality, human-annotated image-text pairs, distributed as 43,906 in the training set, 1,000 in the validation set, and 5,000 in the test set. This data originates from various sources, including: GPT-Synthesized Text Data, Retrieved Image and Text Data, and Diffusion Model-Generated Image Data. Each image-text pair encompasses a varying number of mismatched aspects, ranging from 0 to 3. The collection methods for each data source are described in Section 3.
@@ 3.5 Human Annotation
To standardize the labeling scheme across different data sources and ensure data quality while eliminating potentially harmful content and addressing ethical concerns, we employ a consistent annotation team composed of the same group of workers. The annotation interfaces of different sources of data are shown in Appendix Section 7.
@@ 3.7 Evaluation
where the , and is the weight of the , and , respectively. In this study, we set the weights , , and . To compute the IoU of the aspect representation, we set a threshold to match the predictions with ground truth, if then the predicted triplet matches the ground truth. For each aspect representation prediction, we compute the final score as:
@@
where denotes the number of matched triplets of data , and . We also provide the pseudocode in supplementary material for the ITM-IoU calculation.
@@ 3.9 Qualitative Analysis
We present the analysis of human-annotated and GPT-Synthesized data, focusing on the Vera Score Gap, Grammar Score Gap, and CLIP Score Gap. These metrics compare the quality changes of human-annotated mismatched captions against the original captions. Previous research [14] analyzes the artifact bias through the Vera Score Gap and Grammar Score Gap, and employs the adversarial refinement to fix this bias. In this research, we utilize human annotation to fix this bias and add the CLIP Score Gap to quantify the semantic changes of the human-annotated mismatched captions. The score gap is typically calculated as , where the and are the Vera/Grammar/CLIP scores of the initial captions and the human-annotated GPT-Synthesised data. The findings, illustrated in Figure 6, reveal that score gap distribution lies on the positive spectrum, indicating that hard negative samples in the GPT-Synthesized data exhibit a higher likelihood of being nonsensical. The results also reflect a marginally reduced similarity to the initial matching captions. While these effects are within an acceptable range. Moreover, GPT rewrites do not significantly impact grammar fluency or introduce errors compared to the original captions. We also provide more quality analysis from different perspectives about other data sources in Appendix Section 2.
@@ 5 FineMatch for Text-to-Image Generation Hallucination Detection and Correction
Hallucination issues are prevalent in text-to-image (T2I) generation models, particularly for the text prompts that describe detailed scenes and intricate relationships [43, 28]. Many approaches are proposed for reducing hallucination for T2I models [27, 56, 6, 49, 50, 61]. In this study, we illustrate how FineMatch can help detect and address the mismatch between the text prompts and the generated images for T2I generation models.
@@
The architecture of AutoAlign is shown in Figure 8, as we can see from the diagram that the system is designed to effectively reduce the hallucination for T2I generation. More cases are shown in Appendix 6.